OpenStudy (anonymous):

This question has to do with z-scores... A dean must distribute salary raises to her faculty for next year. She has decided that the mean raise is to be $2000, the standard deviation of raises is to be $400, and the distribution is to be normal. She will attempt to distribute these raises on the basis of merit, meaning that people whose performance is better get better raises. The 5% of the faculty who have done nothing useful in years will receive no more than $ _________ each.

OpenStudy (amistre64):

ahhhh.... standard deviations eh.... where my ti-83 at?

OpenStudy (amistre64):

since it is population; we use mu and such right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but somehow i need to get a z value or an x value for z=x-mean/sd

OpenStudy (amistre64):

2 standards are within the normal right? 95%......

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

z = x-mu/omega omega = sqrt(n.p.q) right? and mu=n.p

OpenStudy (anonymous):

huh? the standard deviation is already given and so is the mean :S

OpenStudy (amistre64):

lol...ok; then mu = 2000 and omega then is 400 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and the 5% is the smaller portion of the distribution. but I don't know how to if its the z score that gives you the 0.05 or if its something else

OpenStudy (amistre64):

use empiraical rule? gotta similar problem here where sd and mean are given

OpenStudy (amistre64):

2000 + 400 is one deviation 2000 + 800 is 2 deviations 2000 + 1200 is 3 deviations right?

OpenStudy (amistre64):

those in the out liers are more than 2 standards so they get a raise of 2000 - 1200 = 800 right?

OpenStudy (amistre64):

make sense?

OpenStudy (anonymous):

yes i think so

OpenStudy (amistre64):

<.......-2 sd ........ -1 sd..........mean .........etc

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

might of gone to far in that but you get the basic premise right?

OpenStudy (anonymous):

lol yea

OpenStudy (amistre64):

good :) we just went over that in stats class yesterday :)

OpenStudy (anonymous):

sweet, so do you have a z table with you?

OpenStudy (amistre64):

i got some stuff in my stats book; that might help

OpenStudy (amistre64):

table a-2; standard deviation (z) distribution

OpenStudy (anonymous):

i have multiple choice answers if it makes a difference. a) 1000, b) 1980, c) 1340 and d)2192 it can't be over 2000 so its one of the first 3 choices :S

OpenStudy (anonymous):

ahh hate stats!

OpenStudy (amistre64):

c looks plausible, but cant really say for sure....

OpenStudy (amistre64):

2000 - 2(400) = 1200; but thats just an educated guess...

OpenStudy (amistre64):

id have to look up how to use a z table to be sure

OpenStudy (anonymous):

well to my understanding, if the 5% is in the smaller portion of the distribution, then you look up0.05 there and get the z score. then just stick that value int he equation and solve for x but its not working

OpenStudy (amistre64):

95% falls at the 2 line; and that result is .4772; or 47.72% at most 2000*.4772 = answer maybe? ....954.4 that dint work

OpenStudy (amistre64):

http://davidmlane.com/hyperstat/z_table.html is a calculator for it i think

OpenStudy (amistre64):

.02275 is what that gives me

OpenStudy (anonymous):

thanks!

OpenStudy (amistre64):

what to do with it I aint got a clue lol

OpenStudy (anonymous):

hmm it won't let me put the 0.5 anywhere though just keeps saying 0.0000

OpenStudy (anonymous):

lol me either.

OpenStudy (amistre64):

i put in the mean 2000 the sd 400 and choose below 1200 for the 5%

OpenStudy (amistre64):

i got a 1342 on the bottom part of it lol

OpenStudy (anonymous):

yea i got that too but i dont really know what it means lol

OpenStudy (amistre64):

i think it means pick c :)

OpenStudy (anonymous):

lol okay. wth! its only one question haha thank you sooooo much for your help!!

OpenStudy (amistre64):

:) hope we are right ;)

OpenStudy (anonymous):

me too :D